OpenStudy (lesteremily7):

Take any right triangle. It can be broken into four similar triangles, each one having edges that are 1/2 the length of their corresponding edges in the big triangle. Use this fact to show that, using the dimension defi nition, any fi lled- in right triangle has dimension 2.

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

What's dimension definition?

OpenStudy (paxpolaris):

fractal dimension... ok

OpenStudy (loser66):

\(D=\dfrac{log4}{log2}=\dfrac{log2^2}{log2}=2\) Dat sit